How are spherical polar coordinates related to the rectangular cartesian coordinates? Illustrate
giving suitable relations. Also write the Schrodinger equation for hydrogen atom in spherical polar coordinates
1
Expert's answer
2020-05-13T14:21:27-0400

Answer:

The spherical polar coordinates "r", "\theta" and "\phi" are related to Cartesian coordinates "x", "y" and "z" via the following equations:

"x = r\text{ sin} \theta\text{ cos} \phi"

"y = r\text{ sin} \theta\text{ sin}\phi"

"z = r \text{ cos}\theta" .

And vice versa:

"r = \sqrt{x^2+y^2+z^2}"

"\phi = \text{arctan} \frac{y}{x}"

"\theta =\text{arctan}\frac{\sqrt{x^2+y^2}}{z}" .

The time-independent Schrödinger equation in spherical polar coordinates for hydrogen atom is:

"\{-\frac{\hbar^2}{2\mu r^2}[\frac{∂}{∂r}(r^2\frac{∂}{∂r}) + \frac{1}{\text{sin}\theta}\frac{∂}{∂\theta}(\text{sin}\theta\frac{∂}{∂\theta}) + \frac{1}{\text{sin}^2\theta}\frac{∂^2}{∂\phi^2}] - \frac{e^2}{4\pi\epsilon_0r}\}\psi(r,\theta,\phi) = E\psi(r,\theta,\phi)"

Solving this equation involves the separation of the variables:

"\Psi(r,\theta,\phi) = R(r)P(\theta)F(\phi)" .